A light bulb rated as 60 W at 220 V has a potential difference of 110 V across its ends. Find the power dissipated in the light bulb, assuming that the light bulb is ohmic (which it likely isn't). Ans: 15 W


1
Expert's answer
2021-12-29T13:06:16-0500



"P=\frac{V^2}{R}"


Where "P=" Power

"V=" Voltage

"R=" Resistance




Resistance of bulb "R=\frac{V^2}{P}=\frac{220^2}{60}=\frac{2420}{3}\Omega"



Power dissipated in the bulb


"P=\frac{110^2}{\frac{2420}{3}}=\frac{110^2}{2420}×3=15W"